# PixelArena: A benchmark for Pixel-Precision Visual Intelligence

### v1

@@ Abstract
Multi-modal large language models that have image output are emerging. Many image generation benchmarks focus on aesthetics instead of fine-grained generation capabilities. In PixelArena, we propose using semantic segmentation tasks to objectively examine their fine-grained generative intelligence with pixel precision. We find the latest Gemini 3 Pro Image has emergent image generation capabilities that generate semantic masks with high fidelity under zero-shot settings, showcasing visual intelligence unseen before and true generalization in new image generation tasks. We further investigate its results, compare them qualitatively and quantitatively with those of other models, and present failure cases. The findings not only signal exciting progress in the field but also provide insights into future research related to multimodality, reasoning, interpretability and benchmarking.
@@ 1 Introduction
Since the release of GPT-4o [gpt4o] in 2024, multi-modal large language models (MLLMs), which have multiple output modalities (e.g., text, images, and audio), have been a focus of research. Numerous models have been developed (e.g., Emu series [emu1, emu2, emu3, emu35], Gemini series [gemini3proimage, gemini25flashimage]). They can generate images based on prompts that include both text and images. This capability is highly malleable, enabling flexible in-context learning and powerful, convenient, and conversational image generation. However, as much focus has been placed on image quality and aesthetics, few have quantitatively examined the precision and generalizability of the image generation capabilities of these models, nor have they examined the limitations of visual reasoning and perception of these models during image generation.
In PixelArena, we propose using pixel-level tasks (e.g., face parsing and general semantic segmentation) to examine MLLMs’ fine-grained control capability and their generalizability in image generation, which we term pixel-precision visual intelligence (PPVI). Specifically, we ask models to perform semantic segmentation (e.g., face parsing) with the CelebAMask-HQ [CelebAMask-HQ] and COCO [coco-dataset] datasets. This allows us to use objective metrics to measure fine-grained generative capability in terms of precision. Qualitatively and quantitatively, we found that Gemini 3 Pro Image [gemini3proimage] represents a significant leap in this front by comparing the results from previous models. With quantitative results, we also show that Gemini 3 Pro Image truly generalizes to new image generation tasks. We also present interesting failure cases and analyze their implications.
@@
    We propose PixelArena, in which pixel-level tasks (e.g., semantic segmentation) are used to quantitatively measure MLLMs’ fine-grained control capability and their generalizability (i.e., PPVI) in image generation tasks.
@@
    We task MLLMs to perform face parsing with the CelebAMask-HQ [CelebAMask-HQ] dataset, revealing surprising emergent zero-shot capabilities in Gemini 3 Pro Image [gemini3proimage]. We also performed experiments to examine potential data contamination in this model, showing that the model did not memorize the answers (i.e., reference masks) but truly understood this image generation task.
@@
    We present the results of semantic segmentation on a more challenging dataset, COCO [coco-dataset], showing that Gemini 3 Pro Image still has reasonable performance and generalization.
  4. 4.
    We have conducted qualitative and quantitative analyzes of the results, including failure cases, hinting at more future directions in multi-modality, reasoning, and interpretability research.
Our website for inspecting the results of PixelArena is available at https://pixelarena.reify.ing.
@@ 2 Related Work
2.1 Semantic Segmentation and Face Parsing
In computer vision research, many segmentation datasets have been developed, such as COCO [coco-dataset], a large-scale benchmark containing object-centric images with pixel-level annotations for diverse everyday scenes; FSS-1000 [fss-1000], a few-shot segmentation dataset featuring 1,000 object categories with only a single annotated example per class; SA-CO [sam3], which extends segment-anything-style annotation to concept-driven segmentation tasks; and SA-MED2D [SA-Med2D], a medical-domain dataset designed to evaluate generalizable 2D segmentation across varied clinical imagery.
We use face parsing with the CelebAMask-HQ [CelebAMask-HQ] dataset as an example, a high-quality facial image dataset that provides detailed pixel-level annotations for 18 distinct facial components across 30,000 celebrity images. Various models have been proposed to push the state of the art on CelebAMask-HQ. The general facial representation model FaRL [general_facial] achieves strong face parsing performance by leveraging large-scale vision–language pretraining and transferring it to CelebAMask-HQ. SegFace [segface] improves the state of the art on CelebAMask-HQ by explicitly addressing long-tail facial components through a balanced segmentation framework.
Another example we present is general semantic segmentation on the panoptic segmentation dataset of the COCO [coco-dataset] dataset. OneFormer [oneformer] and Mask2Former [mask2former] are the state-of-the-art models on this dataset. They are capable of performing universal image segmentation (i.e., semantic segmentation, instance segmentation, and panoptic segmentation) with a single model.
Note that in PixelArena, we do not intend to use MLLMs to compete with these specialized models; instead, we probe the emergent capabilities and generalizability of these generalist models. In our experiments, we task MLLMs to generate masks with published weights or public APIs and no further training.
@@ 2.2 Image Generation Benchmarks and Metrics
Most of the benchmarks [lee2023holisticevaluationtexttoimagemodels, t2ibench, hu2023tifaaccurateinterpretabletexttoimage, yu2022scalingautoregressivemodelscontentrich, saharia2022photorealistictexttoimagediffusionmodels, hu2024ellaequipdiffusionmodels, ku2024imagenhubstandardizingevaluationconditional, zhang2024magicbrushmanuallyannotateddataset, sheynin2023emueditpreciseimage, ye2025imgedit, jayasumana2024rethinkingfidbetterevaluation] for text-to-image generation and image editing focus on evaluating generated natural images rather than the masks that we use. However, due to their diversity and complexity, natural images are difficult to evaluate, and the evaluation metrics are often subject to implicit human preferences or model biases. For example, in HEIM [lee2023holisticevaluationtexttoimagemodels], the metrics used are CLIP score, FID, the score from a LAION aesthetics predictor, human evaluation score, and VQA-based scores. However, the CLIP score, FID, aesthetics score and VQA-based scores may be biased by the models used, while human evaluation is fundamentally based on implicit preferences. These metrics are fundamentally subjective.
In PixelArena, because we task models to generate masks and evaluate the generated masks, we can use standard objective metrics like F1 Score and mIoU.
2.3 Segmentation with MLLMs
Previous research focuses on integrating a MLLM that has multi-modal input but text-only output with a segmentation model. For example, RAS [ras] enhances segmentation models by integrating a mask-centric large multimodal model that selects relevant mask groups from a pool of candidates based on vision-language prompts, enabling flexible and precise mask grouping. SAM4MLLM [sam4mllm] trains a MLLM to output prompts (bounding boxes and points) to guide SAM [sam1] in generating accurate segmentation masks, thus combining language understanding with pixel-level mask generation. In SAM 3 [sam3], the authors propose SAM Agent, which is similar to SAM4MLLM [sam4mllm]. The large multimodal model for remote sensing images [Liu31122025] uses a language model to interpret open-vocabulary queries and conditions a segmentation decoder to produce class-specific masks, enabling flexible, high-resolution semantic segmentation of unseen categories. All of them integrate a MLLM and a segmentation model with text or latent vectors as intermediate representations. However, in PixelArena, we use the original MLLM s without any tools, model extensions, or finetuning. Noteworthy is that in SAM 3, the researcher presented preliminary results generated by Gemini 2.5 Flash Image in object detection tasks with ODinW13 [ODinW13] and RF-100VL [RF-100VL] using prompts and few-shot learning. These results are closely related to ours, but their output is bounding box coordinates, while ours are mask images. We test the finer-grained capabilities of MLLMs at the pixel level with image generation.
3 Methods
@@ 3.1 Dataset
We use subsets of the COCO [coco-dataset] and CelebAMask-HQ [CelebAMask-HQ] datasets as examples.
For the CelebAMask-HQ [CelebAMask-HQ] dataset, we randomly sampled 150 images and their corresponding masks. As the reference masks are $512\times 512$, while the selected MLLMs [gemini25flashimage, gemini3proimage, gpt4o, unimoe2, emu35] natively support image generation with resolutions larger than $512\times 512$ (e.g., $720\times 720$ and $1024\times 1024$), we upsampled the reference masks using nearest neighbors to $1024\times 1024$. We also upsampled the generated masks using nearest neighbors to $1024\times 1024$.
For the COCO [coco-dataset] dataset, we randomly sampled 150 images and their corresponding masks from its panoptic segmentation dataset. We convert the panoptic masks into semantic segmentation masks using its official toolkit [^1] https://github.com/cocodataset/panopticapi. As the resolutions of the images and masks in the dataset are not fixed, we center-crop them based on the shortest dimensions to obtain square ones. Similarly, we upsample the reference masks and generated masks to $1024\times 1024$ using nearest neighbors.
In the following sections, we refer these two subsets as the celeb and coco datasets, respectively. Unless specified otherwise, we evaluate the metrics (i.e., F1 Score, mIoU, and Dice) with upsampled masks.
@@ 3.2 Models and Mask Generation
With different models, we use various methods to generate valid masks.
Figure 1: Palette of the standard color encodings from CelebAMask-HQ [CelebAMask-HQ]
For MLLMs: We select recent models with strong image generation capabilities, including Gemini 3 Pro Image (gmn3) [gemini3proimage], Gemini 2.5 Flash Image (gmn25) [gemini25flashimage], GPT Image 1 (gpti) [gpt4o], Emu 3.5 (emu35) [emu35], and Uni-MoE-2 (unimoe2) [unimoe2]. Note that we tested two variants of unimoe2, Uni-MoE-2 Omni (unimoe2-omni) the flagship model of the series, and Uni-MoE-2 Image (unimoe2-image) the variant finetuned for image generation. For the sake of brevity, we will refer to them by their short code names in the following sections.
As they natively generate images instead of label vectors, we first prompt them to generate images and then convert the pixels from RGB values to class labels. The prompts for the two datasets are composed of three parts: an image from the dataset, an image of the color palette of label encodings (Fig. 1 and 13), and a short text (Listing LABEL:prompt and LABEL:coco-prompt in Supplementary). These prompts provide task specifications and visual grounding for the color encodings, as well as some clarifications. Note that no examples are given, meaning the models have to learn face parsing and general semantic segmentation zero shot. For the sampling parameters of the MLLMs, refer to Table 1 in Supplementary.
Given the mask images from MLLMs, we compare the RGB value of each pixel with the color encodings for the labels, selecting the nearest color and label. It is formulated as Eq. 1, where $\vec{e_{i}}$ is the RGB color vector of label $i$, and $\vec{p}$ is the color vector of a pixel.
$\text{Label Index}\quad i=\operatorname*{argmin}_{i}\ (\vec{e_{i}}\cdot\vec{p})$ (1)
For SAM 3: SAM 3 (sam3) [sam3] accepts text as the prompt for mask generation. We prompt sam3 with the labels of CelebAMask-HQ [CelebAMask-HQ] one by one and merge the corresponding 19 masks into one final mask. For the label of each pixel in the overlapping areas of these masks, we randomly pick one from the overlapping labels.
For specialized computer vision models: We use the pretrained ConvNext [convnext] variant of SegFace (segface) [segface] as a strong baseline model on celeb. On coco, we use OneFormer (1former) [oneformer] as a strong baseline model.
4 Analysis
Due to the stochastic components (e.g., token sampling, diffusion module) in MLLMs, the generation of mask images is inherently stochastic. Therefore, we present results from multiple attempts (i.e., $p=[1,3,5]$).
4.1 Qualitative Comparisons
Figure 2: Comparison between the Results of Different Models on celeb. Not cherrypicked.
Figure 3: Best prediction across celeb by gmn3 with F1 score $0.7030$.
Figure 4: Worst prediction across celeb by gmn3 with F1 score $0.0805$ and parallel attempts.
In Fig. 2, we present results of different models on celeb for qualitative comparison. Among all MLLMs, gmn3 is the only one that understands the task requirements and completes it with high quality. gpti and gmn25 partially understand the task, but gmn25 lacks precise color control or fails to understand the color encodings, while gpti lacks precise control over the composition of the image, hallucinating the upper body of the person. As for sam3, it sometimes misses some labels. emu35 and unimoe2 models completely misundertood the task while they present different failure patterns. emu35 failed to draw plausible masks but we can see that it could control its image generation process to replicate most features in the original image. In contrast, unimoe2 models could not even draw an image similar to the original image, which may be due to its vision system failing to capture the original image, failing to propagate the visual information to its generation module, or failing to control its generation process.
Such failure modes are common in the results of the respective models, implying potential misalignment between the vision system and the generation module or lack of control over the generation process.
We further investigate the results of gmn3. We present the best and worst results from the model in Fig. 3 and Fig. 4. The best prediction has a nearly indistinguishable difference from the reference mask, while the worst prediction is dramatically low in quality. However, the other two attempts in Fig. 4 present reasonable results, which suggests that the generation process is not stable or robust.
Figure 5: Comparison between the reference mask and masks predicted by two strong models.
Even though we do not intend to use MLLMs to compete with the specialized computer vision model (e.g., segface), we present the result in which segface achieved the highest F1 score $0.9539$, while gmn3 achieved $0.6906$ in Fig. 5. We find that these two masks are visually very similar, but the scores have a significant gap, which suggests that we need a new metric for segmentation tasks.
4.2 Quantitative Results and Examining Data Contamination
Figure 6: F1 Scores for experiments on celeb. For $p=[3,5]$, we ask MLLMs to try 3 or 5 times and select the best result in these attempts. As sam3 and segface contain no stochastic components, we did not run more attempts. Due to their poor performance, we did not run experiments of emu35 and unimoe2 with more attempts.
We present the F1 scores in Fig. 6. For mIoU and Dice, please refer to Fig. 14 and Fig. 15 in Supplementary. Aligned with our qualitative analysis, gmn3 achieved the best F1 score, mIoU, and Dice in the selected MLLMs, although it lags behind segface.
As the results of gmn3 are surprisingly good on celeb, we wonder whether data contamination is the cause of such good performance rather than true generalization, since CelebAMask-HQ [CelebAMask-HQ] has published all images and masks on the Internet. We shuffle the color encodings (Fig. 12) instead of using the standard encodings in a new experiment. Noteworthy is that, as shown in Fig. 6, after we shuffled the color encodings, the performance of gmn3 (gmn3-shuffled) did not drop but instead increased by roughly 10% compared to its original result. This means the model did not memorize the reference masks but truly understood the task, including the required color mapping using arbitrary color encodings.
4.3 Further Failure Analysis and Pretended Reflections
We notice that during the image generation process, gmn3 will perform a three-step chain of thoughts (CoT) before presenting the final result: it first considers the task and requirements, then generates a draft image, and finally checks the draft against the requirements and reflects on the result. Such a process seems to imply quality control and iterative refinements. However, as we can see in Fig. 4, a low quality mask could pass its final check in its CoT.
Figure 7: More failure instances. Bottom left: gmn3 correctly identified the left and right eyebrows while confused about the left and right eyes. Bottom right: It mislabeled the hand as cloth.
We further investigate such failures through additional experiments. Although we could not reproduce the extreme failure in Fig. 4, we present two interesting instances in Fig. 7. For the partially incorrect case (bottom left in Fig. 7), gmn3 labeled the right and left eyebrows correctly but mislabeled the eyes, which are in close proximity. In its CoT, however, it explicitly concluded that “I’ve verified that the segmentation mask strictly adheres to all user-specified constraints. Facial feature delineation, including the critical left/right reversal rule, is accurate…” Such pretended quality control also occurred in the second case (bottom right in Fig. 7), in which it misclassified the hand as cloth. In its CoT, it reflected: “… The background, hair, skin, neck, nose, eyebrows, eyes, lips and cloth were all correctly segmented. I’m satisfied that this fulfills all requirements.”
The reflections of the model were merely pretending, blindly affirming the correctness of the result. This may be a fundamental flaw in its multi-modal reasoning. Such examples may also represent adversarial attacks that target potential the flaws and bypass material checking.
4.4 On a more challenging dataset
Even in the failure cases of gmn3 shown in Fig. 4 and 7, we can see that most of them are still reasonably good, which let us wonder whether face parsing on celeb is too easy for advanced MLLMs like gmn3. Therefore, we tested the performance of gmn3, gmn25 and compare them with 1former on coco. Compared to celeb, coco is much more challenging, since its number of classes is much larger (144 vs. 19). We present the best and worst predictions of gmn3 in Fig. 8 and Fig. 9, respectively.
Figure 8: Best prediction of gmn3 on coco, achieving F1 score $0.2695$. The bottom row showcases other four attempts.
Figure 9: Worst prediction of gmn3 on coco, achieving F1 score $0.0$. The buttom row showcases other four attempts.
The best result is interesting in that it seems gmn3 gave up on drawing a detailed mask; yet, the F1 score is higher than other attempts that appear more plausible. Moreover, the issue of pretended reflections persists in this case. The reflection of the model was:
I’m currently verifying the semantic segmentation mask’s consistency with the input image. I’ve analyzed the color mapping for the "bottle" category, expecting a uniform light blue representation. The generated output is a solid light blue square, which accurately reflects the category distribution in the original image. There appear to be no other categories present in the output.
On the other hand, the CoT of the worst result (Fig. 9) reveals a potential fundamental flaw in gmn3’s visual perception module and visual reasoning process. It mistakenly identified the entire picture as “net”, while in other parallel attempts, the vision system could identify details like fences, which implies that its vision system is highly unstable with ragged performance. Furthermore, its multi-modal CoT reasoning could not correct its mistakes due to the issue of pretended reflections. During its examination on the generated mask, it concluded that:
Despite the extreme failures, the outcomes of multiple attempts from gmn3 are still plausibly good, in contrast to the results of gmn25 as shown in Fig. 10 in which gmn25 completely failed to generate a semantic segmentation mask. Our quantitative results in Fig. 11 align with our qualitative analysis as well. For mIoU and Dice, please refer to Fig. 16 and Fig. 17 in Supplementary.
Figure 10: Comparison between the Results of Different Models on coco. We picked the image on which 1former achieved the highest F1 score $0.8096$ while gmn3 achieved $0.1146$.
Figure 11: F1 Scores for experiments on coco. For $p=[3,5]$, we ask MLLMs to try 3 or 5 times and select the best result in these attempts. As 1former contain no stochastic components, we did not run more attempts. Due to their poor performance in previous experiments, we did not run experiments for other MLLMs.
5 Limitations and Discussion
Due to resource constraints, we did not conduct the experiments on the entire CelebAMask-HQ [CelebAMask-HQ] and COCO [coco-dataset] datasets. The quantitative results may be biased towards this subset, which may be easy or difficult. We did not experiment with the models on more segmentation datasets (e.g., SA-CO [sam3]). To our knowledge, SA-CO [sam3] represents one of the hardest segmentation datasets. We would like to test the models with SA-CO [sam3] to understand their performance.
As we can see in Fig. 3 and Fig. 4, one subtle but noticeable difference is in the eye areas. In reference masks, the masks for the eyes cover only the eyeballs, but in the predictions of gmn3, the masks cover most of the periorbital regions. Such a difference may be due to our under-specification of the task, as in our prompt, we did not mention whether it should label only the eyeballs. If we provide instructions that are more detailed, we may be able to improve performance further.
We did not benchmark against RAS [ras], SAM Agent [sam3], or SAM4MLLM [sam4mllm] either. The performance difference between native generation (i.e., ours) and model extension remains to be seen.
For future research, the reason why the shuffled color encodings improve the performance of gmn3 remains an interesting topic that may be closely related to its vision system and visual reasoning capability.
If we could gain access to the source code and weights of the model, the examples of fake quality control may be valuable for mechanistic interpretability research [mechanisticinterpretabilityaisafety] to recover the mechanisms of its internal visual and generative systems.
Regarding benchmark design, we need better metrics than the F1 Score and mIoU to evaluate the performance of MLLMs on semantic segmentation tasks, as we have seen in Fig. 5 that the score discrepancy between segface and gmn3 is significant while it does not reflect the visual similarity between the two masks. Additionally, we may be able to use MLLMs to re-examine the data and labels in existing segmentation benchmarks and refine their data quality, as well as improve the efficiency of annotation in new benchmarks, since many of the results provided by MLLMs are good enough to serve as initial drafts for human annotation or even better than human annotations.
6 Conclusion
We present PixelArena, in which we propose using segmentation tasks to probe the pixel-precision visual intelligence of advanced MLLMs. As two examples, We take face parsing with CelebAMask-HQ [CelebAMask-HQ] and general semantic segmentation with COCO [coco-dataset] to test the PPVI of frontier MLLMs (i.e.[gemini3proimage, gemini25flashimage, gpt4o, emu35, unimoe2, sam3]). We found that Gemini 3 Pro Image represents a major breakthrough in this front. With qualitative and quantitative results, it demonstrates superior performance under our zero-shot setting. We also present failure cases of these models and discuss their failure modes, which may hint at future research directions.
[…]
Supplementary Material
1 Color Palettes and Prompts
Figure 12: Shuffled color palette for CelebAMask-HQ experiments.
We randomly shuffled the color encodings. Fig. 12 shows the shuffled color palette used for the experiments on the CelebAMask-HQ dataset.
Listing. LABEL:prompt is the prompt template we use to prompt MLLMs to generate mask images on the CelebAMask-HQ dataset. As there is ambiguity in the left and right in terms of references (i.e., with respect to images or persons), we clarify this in length to avoid confusion.
When we change the palette for label color encodings, we also change the label encodings in the text prompt.
Listing 1: Prompt Template for our CelebAMask-HQ experiments. We omit the rest of color codings here.
@@ 1 Color Palettes and Prompts
For you convenience, Ive also give you a color palette (the second image) for the label encodings.

### v2

@@ Abstract
Omni-modal models that have multimodal input and output are emerging. However, benchmarking their multimodal generation, especially in image generation, is challenging due to the subtleties of human preferences and model biases. Many image generation benchmarks focus on aesthetics instead of the fine-grained generation capabilities of these models, failing to evaluate their visual intelligence with objective metrics. In PixelArena, we propose using semantic segmentation tasks to objectively examine their fine-grained generative intelligence with pixel precision. With our benchmark and experiments, we find the latest Gemini 3 Pro Image has emergent image generation capabilities that generate semantic masks with high fidelity under zero-shot settings, showcasing visual intelligence unseen before and true generalization in new image generation tasks. We further investigate its results, compare them qualitatively and quantitatively with those of other models, and present failure cases. The findings not only signal exciting progress in the field but also provide insights into future research related to dataset development, omni-modal model development, and the design of metrics.
@@ 1 Introduction
Since the release of GPT-4o OpenAI and et al. (2024) in 2024, omni-modal models (OMMs), which have multiple input and output modalities (e.g., text, images, and audio), have been a focus of research. Numerous OMMs have been developed (e.g., Emu series Sun et al. (2024b, a); Wang et al. (2024); Cui et al. (2025), Gemini series DeepMind (2025b, a)). They can generate images based on prompts that include both text and images. This capability is highly malleable, enabling flexible in-context learning and powerful, convenient, conversational image generation. However, as much focus has been placed on image quality and aesthetics, few have quantitatively examined the precision and generalizability of the image generation capabilities of these models, nor have they examined the limitations of visual reasoning and perception of these models during image generation.
To address the aforementioned issues, in PixelArena, we propose using pixel-level tasks, the ones in semantic segmentation (SS), to examine OMMs’ fine-grained control capability (e.g., painting individual pixels with precise colors) and their generalizability (i.e., generalizing to new pixel-level tasks) in image generation, which we term Pixel-Precision Visual Intelligence (PPVI). By further examining OMMs’ reasoning process, in PixelArena, we can also unveil their limitations in visual reasoning and perception. Specifically, we ask models to perform SS tasks on subsets of CelebAMask-HQ Lee et al. (2020) and COCO Lin et al. (2015) as two examples. This allows us to use objective metrics (e.g., F1 Score, mIoU, and Dice) to measure fine-grained generative capability. We select strong OMMs that were released within the last six months, including Gemini 3 Pro Image DeepMind (2025b), Gemini 2.5 Flash Image DeepMind (2025a), GPT Image 1 OpenAI and et al. (2024), Emu 3.5 Cui et al. (2025), and Uni-MoE-2 Li et al. (2025). In our experiments, we measure quantitative results to evaluate the performance of OMMs on the datasets. We also develop a graphical interface [^1] https://pixelarena.reify.ing to qualitatively examine the results. With these results, we find that Gemini 3 Pro Image represents a significant leap in this front, compared to other models. With the quantitative results, we also show that Gemini 3 Pro Image truly generalizes to new image generation tasks. We also present interesting failure cases and analyze their implications.
@@
    We propose a benchmark, PixelArena, in which pixel-level tasks (i.e., SS tasks) are used to quantitatively measure OMMsPPVI, including fine-grained control capability and generalizability of their image generation capabilities.
@@
    We task OMMs with face parsing using the CelebAMask-HQ Lee et al. (2020) dataset, revealing surprising emergent zero-shot capabilities in Gemini 3 Pro Image DeepMind (2025b). We also perform experiments to examine potential data contamination in this model, showing that the model does not memorize the answers (i.e., reference masks) but truly understands this image generation task. We also present the SS results on a significantly more challenging dataset, COCO Lin et al. (2015), showing that Gemini 3 Pro Image still has reasonable performance and generalization.
@@
    We conduct both qualitative and quantitative analyses of the results, including failure cases, hinting at more future directions in dataset development, OMM research, and design of metrics.
@@ 2 Related Work
2.1 Semantic Segmentation and State-of-the-Art Models
In computer vision research, various segmentation datasets have been developed, such as COCO Lin et al. (2015), a large-scale benchmark containing object-centric images with pixel-level annotations for diverse everyday scenes; FSS-1000 Li et al. (2020), a few-shot segmentation dataset featuring 1,000 object categories with only a single annotated example per class; SA-CO Carion et al. (2025), which extends segment-anything-style annotation to concept-driven segmentation tasks.
In this research, we use face parsing, an SS task, with the CelebAMask-HQ Lee et al. (2020) dataset as an example, a high-quality facial image dataset that provides detailed pixel-level annotations for 18 distinct facial components across 30,000 celebrity images. Another example we present is general SS on the panoptic segmentation dataset of COCO Lin et al. (2015).
Various models have been proposed to push the state of the art on CelebAMask-HQ Lee et al. (2020) and COCO Lin et al. (2015). The latest SegFace Narayan et al. (2025) improves the state of the art on CelebAMask-HQ by explicitly addressing long-tail facial components through a balanced segmentation framework. OneFormer Jain et al. (2023) and Mask2Former Cheng et al. (2022) are the state-of-the-art models on the panoptic segmentation dataset of COCO Lin et al. (2015). They are capable of performing universal image segmentation (i.e., SS, instance segmentation, and panoptic segmentation).
Note that in PixelArena, we do not intend to use OMMs to compete with these state-of-the-art specialized models that are specifically designed and trained for SS on specific datasets; instead, we probe the emergent generative capabilities and generalizability of these generalist models (i.e., OMMs). In our experiments, we task OMMs to generate masks with published weights or public APIs and no further training.
Another line of research focuses on integrating a visual language model (VLM) that has text and image input but text-only output with a segmentation model. RAS Cao et al. (2025) enhances segmentation models by integrating a mask-centric large VLM that selects relevant mask groups from a pool of candidates based on vision-language prompts, enabling flexible and precise mask grouping. SAM4MLLM Chen et al. (2024) trains a VLM to output prompts (bounding boxes and points) to guide SAM Kirillov et al. (2023) in generating accurate segmentation masks, thus combining language understanding with pixel-level mask generation. The large VLM for remote sensing images Liu et al. (2025) uses a language model to interpret open-vocabulary queries and conditions a segmentation decoder to produce class-specific masks, enabling flexible, high-resolution SS of unseen categories. All of them integrate a VLM and a segmentation model with text or latent vectors as intermediate representations. However, in PixelArena, we use the original OMMs without any tools, model integration, or finetuning.
Another noteworthy work is SAM 3Carion et al. (2025). In this work, SAM Agent is proposed, which is similar to SAM4MLLM Chen et al. (2024). It also presents preliminary results generated by Gemini 2.5 Flash Image in object detection tasks with ODinW13 Cappellino et al. (2025) and RF-100VL Robicheaux et al. (2025) using prompts and few-shot learning. This method is similar to ours, but its output is bounding box coordinates, whereas ours are mask images. In contrast to testing visual perception by generating text (i.e., coordinates), we test the finer-grained generative capabilities of OMMs at the pixel level by generating images (i.e., masks). Instead of providing examples, we give high-level instructions to the models to generate masks, forcing them to perform our tasks in zero-shot settings.
@@ 2.2 Image Generation Benchmarks and Metrics
Most of the benchmarks Lee et al. (2023); Huang et al. (2025); Hu et al. (2023); Yu et al. (2022); Saharia et al. (2022); Hu et al. (2024); Ku et al. (2024); Zhang et al. (2024); Sheynin et al. (2023); Ye et al. (2025); Jayasumana et al. (2024) for text-to-image generation and image editing focus on evaluating generated natural images rather than the masks that we use. However, due to their diversity and complexity, natural images are difficult to evaluate, and the evaluation metrics are often subject to implicit human preferences or model biases.
For example, in HEIM Lee et al. (2023), the metrics used are CLIP score, FID, the score from a LAION aesthetics predictor, human evaluation score, and VQA-based scores. However, the CLIP score, FID, aesthetics score and VQA-based scores may be biased by the models used, while human evaluation is fundamentally based on implicit preferences. In ImgEdit Ye et al. (2025), GPT-4o OpenAI and et al. (2024) is prompted with detailed scoring rubrics based on three dimensions (i.e., instruction adherence, image-editing quality, and detail preservation) to score the generated images. It also incorporates a forensic detector, FakeShield Xu et al. (2025), to compute a fake score for the generated images. FakeShield Xu et al. (2025) also uses models for scoring, including GPT-4o OpenAI and et al. (2024) and fine-tuned models based on SAM Kirillov et al. (2023) and Qwen2.5-VL Bai et al. (2025). These metrics are fundamentally subjective with respect to human preferences or model biases.
In PixelArena, because we task models to generate masks and evaluate the generated masks, we can use standard objective metrics such as F1 Score and mIoU.
3 PixelArena
@@ 3.1 Dataset
We use the COCO Lin et al. (2015) and CelebAMask-HQ Lee et al. (2020) datasets as examples. As these datasets contain thousands of images and OMMs are computationally demanding, we randomly sampled 150 images and their corresponding masks from each dataset. With sufficient resources, we can conduct experiments on the entire datasets in the future.
For the CelebAMask-HQ Lee et al. (2020) dataset, we first perform random sampling to obtain a small subset. As the reference masks are $512\times 512$, while the selected OMMs DeepMind (2025a, b); OpenAI and et al. (2024); Li et al. (2025); Cui et al. (2025) natively support image generation with resolutions larger than $512\times 512$ (e.g., $720\times 720$ and $1024\times 1024$), we upsample the reference masks using nearest neighbors to $1024\times 1024$. We also upsample the generated masks using nearest neighbors to $1024\times 1024$.
For the COCO Lin et al. (2015) dataset, we perform the same sampling process on its panoptic segmentation dataset. We convert the panoptic masks into SS masks using its official toolkit [^2] https://github.com/cocodataset/panopticapi. As the resolutions of the images and reference masks in the dataset are not fixed, we center-crop them based on the shortest dimension to obtain square ones. Similarly, we upsample the reference masks and generated masks to $1024\times 1024$ using nearest neighbors. We evaluate the metrics (i.e., F1 Score, mIoU, and Dice) using the processed reference masks and predicted masks generated from the processed images, ensuring the fairness of the evaluation.
In the following sections, we refer to these two subsets as the celeb and coco datasets, respectively. We use three metrics (i.e., F1 Score, mIoU, and Dice) to evaluate the performance of the selected OMMs on the two datasets.
@@ 3.2 Models and Mask Generation
For different models, we use different methods to generate valid segmentation masks. For the sake of brevity, we will refer to the selected models by their short code names (e.g., gmn3) in the following sections.
Figure 1: Palette of the standard color encodings from CelebAMask-HQ Lee et al. (2020).
@@
For your convenience, Ive also given you a color palette (the second image) for the label encodings.